OpenStudy (anonymous):

Does this integral have a closed form? \[\newcommand{arctanh}{\mathbin{\text{arctanh}}} \int_{-1}^1\frac{\arctan x}{\arctanh\, x}\,dx\]

OpenStudy (anonymous):

(Quick note: I've defined `\arctanh` in LaTeX, so anyone in this thread should be able to use it just fine)

OpenStudy (anonymous):

Here's a preliminary numerical result due to Mathematica

OpenStudy (kainui):

I want to say yes cause I've never heard of an integral being multivalued before but I'm suspicious haha. http://www.wolframalpha.com/input/?i=integral+from+-1+to+1+arctan+x+%2F+arctanh+x++dx

OpenStudy (anonymous):

What I meant by "exact value" was "closed form" :)

ganeshie8 (ganeshie8):

\[\large \int_{-1}^1\frac{\arctan x}{\arctanh\, x}\,dx = 2\int\limits_{-\pi/4}^{\pi/4}\dfrac{u\sec^2u}{\ln|\frac{1+\tan u}{1-\tan u}|}du\]

OpenStudy (kainui):

Ohhhh ok, sorry it's still early for me here haha.

OpenStudy (kainui):

Is there a way to calculate an integral of a ratio as the ratio of the integrals or something like that?

ganeshie8 (ganeshie8):

\[\large \int_{-1}^1\frac{\arctan x}{\arctanh\, x}\,dx = 2\int\limits_{-\pi/4}^{\pi/4}\dfrac{u\sec^2u}{\ln|\frac{1+\tan u}{1-\tan u}|}du = 4\int\limits_{0}^{\pi/4}\dfrac{u\sec^2u}{\ln|\frac{1+\tan u}{1-\tan u}|}du\] idk if it can be worked using simple definite integral props, but im giving it a try first

OpenStudy (kainui):

I wonder if there's a geometric way to solve this integral

OpenStudy (kainui):

Here's an idea, can we make this substitution does it help us? \[-i \tan^{-1}(i x) = \tanh^{-1}(x)\]

OpenStudy (kainui):

Or perhaps rewriting as: \[2i \int\limits_0^1 \frac{\ln \frac{1-ix}{1+ix}}{\ln \frac{1+x}{1-x}} dx = 2i \int\limits_0^1 \log_{\frac{1+x}{1-x}} \left( \frac{1-ix}{1+ix} \right) dx\]

OpenStudy (kainui):

Ok ok ok how about from here can we turn this into a double integral perhaps? \[2i \int\limits_0^1 \frac{\ln \frac{1-ix}{1+ix}}{\ln \frac{1+x}{1-x}} dx\] I want to make the substitution \[1+ix = r e^{i \theta}\]\[ idx = e^{i\theta} dr + ir e^{i \theta} d \theta \] and looking at this \[1+ix = r e^{i \theta}\] I change the bounds from \[0 \le x \le 1\] to \[0 \le \theta \le \frac{\pi}{2}\]\[1 \le r \le \sqrt{2} \] And then plug this in somehow I guess? Just an idea.

OpenStudy (kainui):